OpenStudy (lifeisadangerousgame):

Can someone help me understand builder notation in Algebra? I can't understand the concept at all...

terenzreignz (terenzreignz):

set-builder notation?

OpenStudy (lifeisadangerousgame):

Yes

terenzreignz (terenzreignz):

Okay, let's have a look at a rather easy example... This set: \[\Large \{2,4,6,8,10...\}\] This is the roster form of the set of positive even numbers, right?

OpenStudy (anonymous):

Well,I think u must watch this

OpenStudy (anonymous):

https://www.youtube.com/watch?v=xnfUZ-NTsCE

OpenStudy (lifeisadangerousgame):

Right*

OpenStudy (anonymous):

@LifeIsADangerousGame https://www.youtube.com/watch?v=xnfUZ-NTsCE

OpenStudy (lifeisadangerousgame):

kk, I'll watch that video Udit

terenzreignz (terenzreignz):

Nice video.

OpenStudy (anonymous):

Well, yes

OpenStudy (lifeisadangerousgame):

Okay, it makes more sense after watching the video, Is x the numbers between the brackets?

terenzreignz (terenzreignz):

Yes... the numbers between the curly brackets should be all the values x could possibly take. And that set we have above was the second set the guy handled... he defined it as \[\Large \{2x \ \left| \ x\in \mathbb{N} \right.\}\]

OpenStudy (anonymous):

Well do u know khanacademy

terenzreignz (terenzreignz):

Personally, I prefer \[\Large \{x \ \left| \ x = 2k \ ; \ k\in \mathbb{N} \right.\}\]

OpenStudy (lifeisadangerousgame):

So the numbers is the brackets is what x could be, not what x is going to be if you take one of the numbers from inside of the brackets and plug it into x?

terenzreignz (terenzreignz):

Think of it this way.... \[\Large \{x \ \left| \ \color{red}{\text{an appropriate description so that no value is missed}} \right. \}\]

terenzreignz (terenzreignz):

Here's a trickier one: It's related to his last example: \[\Large \{2,3,5,6,7,8,10,11, 12...\}\]

terenzreignz (terenzreignz):

Basically, we're listing all the positive integers (natural numbers) which are NOT perfect squares... how do we proceed in writing this in set-builder?

terenzreignz (terenzreignz):

So... let's have a little step-by-step; \[\Large \color{red}{\{x \ |}\] This means "the set of all x such that..."

terenzreignz (terenzreignz):

Okay, so now, we describe ALL the elements of that set... how would you describe them?

OpenStudy (lifeisadangerousgame):

its plus 1 for a few numbers then plus 2 and so on?

terenzreignz (terenzreignz):

Well, not like that... something that's more general... they're all positive integers (natural numbers) aren't they?

OpenStudy (lifeisadangerousgame):

They're all natural numbers and they go on until infinity?

terenzreignz (terenzreignz):

"They're all natural numbers" is enough... let's put that down... \[\Large \color{red}{\{x \ \left| \ \color{green}{x\in \mathbb{N} }\right. } \]

terenzreignz (terenzreignz):

Are we done here?

OpenStudy (lifeisadangerousgame):

I don't think so, its all natural numbers, but there's a patten I think

terenzreignz (terenzreignz):

The answer is no...we can't end it here quite yet... if we do, the set would erroneously include ALL natural numbers, including 1, 4, 9, 16 etc... which aren't supposed to be in the set...

terenzreignz (terenzreignz):

Just a reminder: this: \[\Large \{2,3,5,6,7,8,10,11, 12...\}\] is our set... set of all natural numbers which aren't perfect squares...

terenzreignz (terenzreignz):

Maths is like a language itself... how do you say "x is not a perfect square" in mathematical-symbols?

OpenStudy (lifeisadangerousgame):

x^1?

terenzreignz (terenzreignz):

Much like a language, there are 'sentences' and 'phrases' in maths... and that's just a phrase :D "x is not a perfect square" is a sentence... it must contain an equation or an inequality to be a sentence :D

terenzreignz (terenzreignz):

This is a phrase: \[x + y\] (sum of x and y) And this is a sentence: \[x +y= z\](The sum of x and y is z.)

terenzreignz (terenzreignz):

Okay, to say that x is not a perfect square, we can write... \[\Large \color{red}{\{x \ \left| \ \color{green}{x\in \mathbb{N} \ ; \color{blue}{ \ x \ne k^2} }\right. }\] Is that enough?

OpenStudy (lifeisadangerousgame):

No we need to do 2 < x < oo (infinity) I think?

terenzreignz (terenzreignz):

The fact that we have \(\large x \in \mathbb{N}\) already says that x goes to infinity... it's important not to be redundant...

OpenStudy (lifeisadangerousgame):

Oh then maybe just 2 < x?

terenzreignz (terenzreignz):

Not quite... there's only one more thing to add... keep in mind that we just added a new variable, k. We now have to define what k is. In particular, x is not a perfect square if x is not equal to k^2 for any integer k, right?

terenzreignz (terenzreignz):

So finally, we add this detail: \[\Large \color{red}{\{x \ \left| \ \color{green}{x\in \mathbb{N} \ ; \color{blue}{ \ x \ne k^2} } \ ; \ k \in \mathbb{N}\right. } \ \}\] And we're done...

terenzreignz (terenzreignz):

So we read this as: The set of all x such that x is a natural number where x is not equal to k^2 for any natural number k. Indeed, that gives us all the positive integers which are not perfect squares.

terenzreignz (terenzreignz):

By the way, there is more than one way to say that... there is something far simpler :D But (in my opinion) trickier to spot... one can also have: \[\Large \{x \ \left| \ \color{blue}{\sqrt x \notin \mathbb{N}}\right.\}\]

terenzreignz (terenzreignz):

Or the set of all x where the square root of x is NOT an integer.

terenzreignz (terenzreignz):

Wait, sorry... \[\Large \{x \ \left| \ \color{green}{x\in \mathbb{N} } \ \color{blue}{ \ ; \sqrt x \notin \mathbb{N}}\right.\}\]

terenzreignz (terenzreignz):

^ That was simpler... For most sets, there is more than one way to write it in set-builder notation... what's important is that they all refer to one set.

terenzreignz (terenzreignz):

Don't worry if that confused you... that was a confusing set to begin with... Can you try this: \[\Large \{1,3,5,7,9...\}\]? Clearly the set of all positive odd numbers... how would you write this in set-builder?

OpenStudy (lifeisadangerousgame):

I apologize for not answering, I felt sick and had to brb. I'll try to problem now